Determine the molar solubility of lead (II) iodide, PbI2, in 0.0500 mol/L NaI.
Answer #1

Let x = mol/L of PbI2 that dissolve
This will give us x mol/L Pb2+ and 2x mol/L I-

Ksp of PbI2 = 8.40 x 10^-9

we have arleady some I- : its concentration is 0.0500 M

at equilibrium : [Pb2+]= x and [I-] = 2x + 0.0500

Ksp = [Pb2+][I-]^2 = (x)(2x + 0.0500)^2

x = molar solubility = 3.36 x 10^-6 M
Answer #3

Let x = mol/L of PbI2 that dissolve

This will give us x mol/L Pb2+ and 2x mol/L I-


Ksp of PbI2 = 8.40 x 10^-9


we have arleady some I- : its concentration is 0.0500 M


at equilibrium : [Pb2+]= x and [I-] = 2x + 0.0500


Ksp = [Pb2+][I-]^2 = (x)(2x + 0.0500)^2


x = molar solubility = 3.36 x 10^-6 M